OpenStudy (anonymous):

I need help on how to answer this problem and others like it: 5/(x-1)=3/(x+2)

OpenStudy (anonymous):

try to time both sides with (x-1)(x+2)

OpenStudy (anonymous):

I've done that but I don't now how to procede from there

OpenStudy (anonymous):

can you write it here what result do you get after you've done that?

OpenStudy (anonymous):

5(x+2)=3(x-1)

OpenStudy (anonymous):

then expand the brackets, what u get?

OpenStudy (anonymous):

5x+10=3x+3

OpenStudy (anonymous):

subtract both sides for 3x?

OpenStudy (anonymous):

Wow I got it from here thanks alot

OpenStudy (anonymous):

no problem